OpenStudy (anonymous):

what is the derivative of y = 4throot(x)

OpenStudy (anonymous):

d(x^n)/dx = n ( x)^ [n-1] for 4th root x , n= 1/4 so just calculate now. :)

OpenStudy (anonymous):

\[\sqrt[4]{x}=x^{\frac{1}{4}}\] the power rule

OpenStudy (anonymous):

so the derivative is 1/4(x)^(-3/4) ?

OpenStudy (anonymous):

yes.

OpenStudy (amistre64):

ive never seen a derivative for a "throot" before ...

OpenStudy (anonymous):

sore subject

OpenStudy (anonymous):

you might want to make sure you know that this is \[\frac{1}{4\sqrt[4]{x^3}}\]